# LagCeresa.bib

### v1

@@ 1. Introduction
In a smooth projective variety, studying and comparing the different equivalence relations between algebraic cycles can shed light on a wealth of interesting geometric phenonema. A chief example is that of the Ceresa cycle of a curve, which is a $1$-cycle living its Jacobian, and provided the first instance of a cycle which was shown to be homologically, but not algebraically, trivial [Ceresa]. This discrepancy between homological and algebraic equivalence is encoded in the Griffiths group of the variety. Our aim is to propose a symplectic counterpart to this group and use it to exhibit, through an $SYZ$ picture, a symplectic manifestation of the fact that the Ceresa cycle is not algebraically trivial.
While we survey below how the (cylindrical) cobordism group of a symplectic manifold $(X,\omega)$ is known to be closely related to the Chow groups of its mirror, a candidate mirror to its Griffiths groups (that is, the quotient of the Chow groups by algebraic equivalence) has, to the best of our knowledge, not yet appeared in the literature. We propose such a mirror by introducing an equivalence relation which we call algebraic Lagrangian cobordism on the set of suitable Lagrangians in a symplectic manifold, and from this we define the algebraic Lagrangian cobordism group of $(X,\omega)$, $algCob(X)$.
The predicted relation between the Lagrangian cobordism group of a symplectic manifold $X$, denoted $\mathcal{C}ob(X)$, and the Chow groups of its mirror $\check{X}$, was first brought to light in the context of SYZ fibrations by Sheridan–Smith in [nick&ivan1]. Their constructions rely on the following observation. Let $B$ be a tropical torus, $Z_{1}$ and $Z_{2}$ be tropical cycles in $B$, and consider the Lagrangian torus fibration
@@
$\mathcal{D}^{\pi}\mathcal{F}uk(X)\cong D^{b}Coh(Y)$
@@
where $CH_{*}(Y)$ is the Chow ring of $Y$ and the second isomorphism is given by the Chern character map $ch$; see [gillet_Ktheory_intersection_theory, Section 2.3]. On the left-hand side, Biran--Cornea have shown in [bc1, Biran_Cornea_2] that $K_{0}(D\mathcal{F}uk(X))$ is closely related to (the appropriate flavour of) the Lagrangian cobordism group $\mathcal{C}ob(X)$, which we define in section 2. More precisely, Lagrangian cobordisms yield triangle decompositions in the Fukaya category, therefore there is a surjective morphism
$\mathcal{C}ob(X)\longrightarrow K_{0}(\mathcal{F}uk(X)).$
@@
In some cases, we know this to be an isomorphism [haug_T2, alvaro_bielliptic, unob_lag_cob_gps_surfaces], and it is interesting to note that the first two of these examples rely on a proof of homological mirror symmetry hence on understanding the mirror Chow groups. Purely symplectically, this line of questions amounts to uncovering which triangle decompositions in $\mathcal{F}uk(X)$ can be realised geometrically by Lagrangian cobordisms.
@@
For symplectic topologists, one of the motivations for studying the groups $\mathcal{C}ob(X)$ associated to a symplectic manifold $X$ is that Lagrangian cobordism provides a coarser equivalence relation than Hamiltonian isotopy (see Example 2.2). Therefore understanding $\mathcal{C}ob(X)$ ought to be simpler than the famously hard problem of classifying Lagrangians in $X$ up to Hamiltonian isotopy. Nevertheless, currently there are only few instances in which these cobordism groups are fully understood. In dimension two, Arnold computed $\mathcal{C}ob(T^{*}S^{1})$ using flux arguments [Arnold_Lag&Leg_cobordisms] (see Section 2), then Haug used similar techniques to compute $\mathcal{C}ob(T^{2})$ [haug_T2]. Rathael-Fournier computed Lagrangian cobordisms groups for higher-genus surfaces through a different approach [unob_lag_cob_gps_surfaces], namely using the action of the mapping class group to obtain generators following work of Abouzaid [abouzaid_fukaya_cat_higher_genus]. The first computation for compact symplectic manifolds in dimension four was carried out by Muñiz-Brea in [alvaro_bielliptic] for symplectic bielliptic surfaces, restricting to tropical Lagrangians. In the non-compact case, Bosshard studies $\mathcal{C}ob(X)$ when $X$ is a Liouville manifold, and computes these for non-compact Riemann surfaces of finite type [bosshard_Lag_cob_Liouville]. Here again, we have omitted information about the exact flavour of cobordism group considered in each of these cases.
In light of the observations above, it is natural to wonder whether one could further our understanding of Lagrangian cobordism groups by translating interesting examples about Chow cycles to the symplectic realm. Our main result achieves this; we exhibit an example of a nullhomologous Lagrangian $3$-manifold, namely the Lagrangian Ceresa cycle inside a symplectic $6$-torus, which is non-zero (and even non-torsion) in its Lagrangian cobordism group.
@@
The Lagrangian Ceresa cycle $L-L^{-}$ has infinite order in $\mathcal{C}ob^{or}(X)$.
Here $X$ is a $6$-torus endowed with a symplectic structure constructed in Section 6; see Section 7 for a more precise version of this statement.
This result mirrors a classical algebraic story of the Ceresa cycle. Given an algebraic curve $C$, its Ceresa cycle is a homologically trivial $1$-cycle in the Jacobian $J(C)$, which we denote by $C-C^{-}\in CH_{1}(J(C))_{hom}$. It has been the subject of extensive study since Ceresa proved that for a generic genus $3$ curve, it provides an example of a homologically trivial cycle which is not algebraically trivial. More precisely, letting
$\text{Griff}_{1}(J(C)):=CH_{1}(J(C))_{hom}/\sim$
be the Griffiths group, with $\sim$ denoting algebraic equivalence, Ceresa proves the following:
Theorem 1.2 ([Ceresa]).
@@
For a generic genus $3$ curve, $C-C^{-}$ has infinite order in $\text{Griff}_{1}(J(C))$.
Let us note that Theorem 1.1 is not exactly mirror to Ceresa’s result 1.2, but rather to the weaker statement that for a generic genus $3$ curve, $C-C^{-}$ has infinite order in $CH_{1}(J(C))_{hom}$. This is where algebraic Lagrangian cobordisms come into play. The reason that the equivalence relation mirror to algebraic - rather than rational - equivalence is less straightforward, is that mirrors to higher genus curves are vastly complicated objects. An insight by which one can circumvent this difficulty is that a curve can be embedded into its Jacobian. From this perspective one can equivalently define algebraically equivalent cycles as cycles living in some family over a abelian variety. Then a candidate mirror equivalence relation for Lagrangians in $(X,\omega)$ would be to say two Lagrangians $L$ and $L^{\prime}$ are algebraically Lagrangian cobordant if they live in some Lagrangian family over a symplectic polarised torus; see Section 3. We call the group of suitable Lagrangians modulo the equivalence relation generated by algebraic Lagrangian cobordisms the algebraic Lagrangian cobordism group of the symplectic manifold, which we denote $algCob(X,\omega)$ or simply $algCob(X)$. A Lagrangian cobordism between finite sets of Lagrangians implies in particuar that they are algebraically Lagrangian cobordant, so there is a natural projection map
$\mathcal{C}ob(X)\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}algCob(X).$
With this set up, Theorem 1.1 can be upgraded to
@@
The Lagrangian Ceresa cycle $L-L^{-}$ has infinite order in $algCob^{or}(X)$.
Our proof relies on the tropical version of Ceresa’s statement 1.2, which is due to Zharkov [zharkov_tropical_ceresa]. To a tropical curve $C$, one can similarly associate its tropical Ceresa cycle $C-C^{-}\in CH_{1}(J(C))^{trop}_{hom}$. Zharkov proves

### v2

# Algebraic Lagrangian cobordisms, flux and the Lagrangian Ceresa cycle

@@ 1. Introduction
In a smooth projective variety, studying and comparing the different equivalence relations between algebraic cycles can shed light on a wealth of interesting geometric phenonema. A chief example is that of the Ceresa cycle of a curve, which is a $1$-cycle living in its Jacobian, and provided one of the first instances of a cycle which was shown to be homologically but not algebraically trivial [Ceresa]. This discrepancy between homological and algebraic equivalence is encoded in the Griffiths group of the variety. Our aim is to propose a symplectic counterpart to this group and use it to exhibit, through an $SYZ$ picture, a symplectic manifestation of the fact that the Ceresa cycle is not algebraically trivial.
We survey below how the (cylindrical) Lagrangian cobordism group of a symplectic manifold $(M,\omega)$ is known to be closely related to the Chow groups of its mirror. However, a candidate mirror to its Griffiths groups (that is, the quotient of the Chow groups by algebraic equivalence) has, to the best of our knowledge, not yet appeared in the literature. We propose such a mirror by introducing an equivalence relation which we call algebraic Lagrangian cobordism on the set of suitable Lagrangians in a symplectic manifold, and from this we define the algebraic Lagrangian cobordism group of $(M,\omega)$, $algCob(M,\omega)$.
While rational equivalence is defined in terms of flat families of cycles over $\mathbb{P}^{1}$, algebraic equivalence is defined in terms of flat families of cycles over curves of arbitrary genus. These are known to have vastly more complicated mirrors, making a translation into symplectic geometry more subtle. The insight we use to circumvent this difficulty is that higher genus curves can be embedded in their Jacobian torus, which are abelian varieties. These have SYZ mirrors which we call polarised symplectic tori. From here, we define algebraic Lagrangian cobordisms in terms of “Lagrangian families” over a polarised symplectic torus. More explicitly, an algebraic Lagrangian cobordism between two Lagrangians $L$ and $L^{\prime}$ in a symplectic manifold $(M,\omega)$ is a Lagrangian correspondence $\Gamma\subset(X(B)\times M,-\omega_{X(B)}\oplus\omega)$ such that $\Gamma(F_{p})=L$ and $\Gamma(F_{q})=L^{\prime}$, where $(X(B),\omega_{X(B)})$ is a symplectic polarised torus and $F_{p},F_{q}\subset X(B)$ are two Lagrangian tori (Definition 4.3). The algebraic Lagrangian cobordism group of $(M,\omega)$, $algCob(M,\omega)$, is then defined as the quotient of integer combinations of Lagrangians by the equivalence relation generated by algebraic Lagrangian cobordism.
The predicted relation between the Lagrangian cobordism group of a symplectic manifold $M$, denoted $\mathcal{C}ob(M,\omega)$, and the Chow groups of its mirror $Y$, was first brought to light in the context of SYZ fibrations by Sheridan–Smith in [nick_ivan1]. Their constructions rely on the following observation. Let $B$ be a tropical torus, $Z_{1}$ and $Z_{2}$ be tropical cycles in $B$, and consider the Lagrangian torus fibration
@@
$\mathcal{D}^{\pi}\mathcal{F}uk(M)\cong D^{b}Coh(Y)$
@@
where $CH_{*}(Y)$ is the Chow ring of $Y$ and the second isomorphism is given by the Chern character map $ch$; see [gillet_Ktheory_intersection_theory, Section 2.3]. On the left-hand side, Biran-Cornea have shown in [bc1, Biran_Cornea_2] that $K_{0}(D\mathcal{F}uk(M))$ is closely related to (the appropriate flavour of) the Lagrangian cobordism group $\mathcal{C}ob(M)$, which we define in section 3. More precisely, Lagrangian cobordisms yield triangle decompositions in the Fukaya category, therefore there is a surjective morphism
$\mathcal{C}ob(M)\longrightarrow K_{0}(\mathcal{F}uk(M)).$
@@
In some cases, we know this to be an isomorphism [haug_T2, alvaro_bielliptic, unob_lag_cob_gps_surfaces], and it is interesting to note that the first two of these examples rely on a proof of homological mirror symmetry hence on understanding the mirror Chow groups. Purely symplectically, this line of questions amounts to uncovering which triangle decompositions in $\mathcal{F}uk(M)$ can be realised geometrically by Lagrangian cobordisms.
@@
For symplectic topologists, one of the motivations for studying the groups $\mathcal{C}ob(M)$ associated to a symplectic manifold $M$ is that Lagrangian cobordism provides a coarser equivalence relation than Hamiltonian isotopy (see Example 3.2). Therefore understanding $\mathcal{C}ob(M)$ ought to be simpler than the famously hard problem of classifying Lagrangians in $(M,\omega)$ up to Hamiltonian isotopy. Nevertheless, currently there are only few instances in which these cobordism groups are fully understood. In dimension two, Arnold computed $\mathcal{C}ob(T^{*}S^{1})$ using flux arguments [Arnold_Lag_Leg_cobordisms] (see Section 3), then Haug used similar techniques to compute $\mathcal{C}ob(T^{2})$ [haug_T2]. Rathael-Fournier computed Lagrangian cobordisms groups for higher-genus surfaces through a different approach [unob_lag_cob_gps_surfaces], namely using the action of the mapping class group to obtain generators following work of Abouzaid [abouzaid_fukaya_cat_higher_genus]. The first computation for compact symplectic manifolds in dimension four was carried out by Muñiz-Brea in [alvaro_bielliptic] for symplectic bielliptic surfaces, restricting to tropical Lagrangians. In the non-compact case, Bosshard studies $\mathcal{C}ob(M)$ when $M$ is a Liouville manifold, and computes these for non-compact Riemann surfaces of finite type [bosshard_Lag_cob_Liouville]. Here again, we have omitted information about the exact flavour of cobordism group considered in each of these cases.
In light of the observations above, it is natural to wonder whether one could further our understanding of Lagrangian cobordism groups by translating interesting facts about Chow groups to the symplectic realm. More ambitiously, after having introduced algebraic Lagrangian cobordisms as an equivalence relation on Lagrangians “mirror” to algebraic equivalence of cycles, one could wonder whether information about Griffiths groups can also be accessed symplectically. With these objectives in mind, the story of the Ceresa cycle is a natural candidate to turn one’s attention to.
Our main Theorem builds upon a tropical version of the Ceresa cycle story, worked out by Zharkov in [zharkov_tropical_ceresa]. To a certain type of genus $3$ tropical curve $C$, we associate a nullhomologous Lagrangian $3$-manifold $L_{C}-L_{C}^{-}$ called the Lagrangian Ceresa cycle of $C$. Its ambient symplectic manifold is a symplectic $6$-torus $X(J(C))$, also built out of $C$. We prove the following:
@@
For generic tropical curves $C$ of a certain type, the Lagrangian Ceresa cycle $L_{C}-L_{C}^{-}$ has infinite order in $algCob^{or}(X(J(C)))$.
Here $algCob^{or}(X(J(C)))$ is the oriented algebraic Lagrangian cobordism group of $X(J(C))$.
Our constructions for $L_{C}-L_{C}^{-}$ as well as $X(J(C))$ build upon those of Zharkov, as does the proof of 1.1. Therefore we begin with an outline of the tropical story. More details, as well as the relevant background in tropical geometry can be found in Section 2.1.
In [zharkov_tropical_ceresa], Zharkov considers a genus $3$ tropical curve $C$ of type $K4$, represented below in Figure 1. Its data is that of the underlying graph, and the choice of real positive lengths $a,b,c,d,e,f$ for the edges $A,B,C,D,E,F$.
ppEFDACB
Figure 1. A tropical curve of type $K4$.
After an arbitrary choice of basepoint, $C$ can be embedded in its tropical Jacobian $J(C)$, which in this case is a $3$-dimensional tropical torus. The image of this embedding is represented in Figure 4. This Jacobian has a natural group structure, therefore an involution by taking inverses for the group law. The Ceresa cycle of $C$ is a nullhomologous $1$-cycle in $J(C)$: it is the difference between the image of the curve in its Jacobian, and its image composed with this involution. We write it as $C-C^{-}\in\text{Griff}^{trop}_{1}(J(C))$, where $\text{Griff}^{trop}_{1}(J(C))$ denotes the $1$-dimensional tropical Griffiths group of $J(C)$. Zharkov proves the following:
Theorem 1.2 ([zharkov_tropical_ceresa, Ceresa]).
@@
For a generic genus $3$ tropical curve of type $K4$, $C-C^{-}$ is not trivial in $\text{Griff}^{trop}_{1}(J(C))$.
This is a tropical version of a famous Theorem of Ceresa [Ceresa].
The word generic here refers to the choice of edge lengths of $C$, which are $6$ positive real numbers, being generic. The proof of Theorem 1.2 is sketched as follows. Zharkov shows that if two $1$-cycles in $J(C)$ are algebraically equivalent (in particular they are homologous), then certain integrals over $2$-chains realising their homological equivalence are constrained. More precisely, he defines a determinantal form $\Omega_{0}$: this is a $2$-form with coefficients encoding information about the tropical structure of $C$. He shows that if $C$ and $C^{-}$ were algebraically equivalent in $J(C)$, then for any $2$-chain $\gamma$ with $\partial\gamma=C-C^{-}$ we would have
$\int_{\gamma}\Omega_{0}\in\mathcal{P}(\Omega_{0}),$
where $\mathcal{P}(\Omega_{0})$ is the set of periods of $\Omega_{0}$. He then constructs an explicit such $2$-chain $\gamma_{0}$ and computes
$\int_{\gamma_{0}}\Omega_{0}=-ad.$
After a straightforward verification, one observes that ad-ad is generically not in $\mathcal{P}(\Omega_{0})$, and concludes.
We now return to the statement of Theorem 1.1. To the tropical torus $J(C)$, one can associate a symplectic manifold
$X(B):=T^{*}J(C)/T^{*}_{\mathbb{Z}}J(C),$
where $T^{*}_{\mathbb{Z}}J(C)\subset T^{*}J(C)$ is a full rank lattice determining the tropical structure of $J(C)$. This is a $6$-dimensional symplectic torus constructed solely from the data of the tropical curve $C$. In fact, the cohomology class of the symplectic form is determined by its nine non-vanishing periods, which are the entries of the polarisation matrix of the curve $C$:
$Q=\begin{pmatrix}a+e+f&-f&-e\\ -f&b+d+f&-d\\ -e&-d&c+d+e\end{pmatrix}$.
The Lagrangian Ceresa cycle $L_{C}-L_{C}^{-}$ associated to $C$ is the Lagrangian lift of $C-C^{-}$ to the Lagrangian torus fibration
$X(J(C))\longrightarrow J(C).$
Each of its two embedded pieces is an oriented graph $3$-manifold built out of the $K4$ graph from Figure 1, with first and second Betti numbers $6$.
The proof of Theorem 1.1 can be sketched as follows. We show in Proposition 4.3 that if $L_{C}$ and $L_{C}^{-}$ were algebraic Lagrangian cobordant (in particular homologous), then for any $4$-chain $\gamma$ in $X(J(C))$ with $\partial\gamma=L_{C}-L_{C}^{-}$, we would have
$\int_{\gamma}\omega^{2}\in\mathcal{P}(\omega^{2}),$
where $\mathcal{P}(\omega^{2})$ denotes the set of periods of $\omega^{2}$.
Because $L_{C}-L_{C}^{-}$ is a Lagrangian lift to $X(J(C))$ of $C-C^{-}\subset J(C)$, we can build a $4$-chain $\hat{\phi}(\gamma_{0})$ in $X(J(C))$ satisfying $\partial\hat{\phi}(\gamma_{0})=L_{C}-L_{C}^{-}$ by “lifting” Zharkov’s $2$-chain $\gamma_{0}$ in $J(C)$. This involves a conormal-type construction which has the virtue of satisfying
(1.1) $\int_{\hat{\phi}(\gamma_{0})}\omega^{2}=\int_{\gamma_{0}}\Omega_{0}.$
This equality stems from the fact that the $4$-form $\omega^{2}$ on $X(J(C))$ can be obtained from the $2$-form $\Omega_{0}$ on $J(C)$ through a construction “dual” to the one taking $\gamma_{0}$ to $\hat{\phi}(\gamma_{0})$. In particular, we have
$\mathcal{P}(\omega^{2})=\mathcal{P}(\Omega_{0}).$
Putting this together with Zharkov’s computation yields
$\int_{\hat{\phi}(\gamma_{0})}\omega^{2}=-ad\notin\mathcal{P}(\omega^{2}),$
from which we conclude.
In is important to emphasize that the theory we build in this paper is much more general than what is strictly necessary for this proof. Equation (1.1) above is merely an instance of a much more general correspondence we display between what we call tropical fluxes and symplectic fluxes.
Tropical fluxes are morphisms
$\Theta^{trop}:\text{Griff}_{k}^{trop}(B)_{hom}\longrightarrow\mathbb{R}/\mathcal{P}_{k},$
where $B$ is any tropical torus, $\mathcal{P}_{k}\subset\mathbb{R}$ are finitely generated groups of periods, and $\text{Griff}_{k}^{trop}(B)_{hom}$ denotes the homologically trivial part of the tropical Griffiths group of $B$. These morphisms are given by integrals over $(k+1)$-chains bounding the homologically trivial cycles. They are constructed in Section 2.2, using the language of tropical (co)homology.
Similarly, symplectic fluxes are morphisms
$\Theta^{symp}:algCob^{or}(M,\omega)_{hom}\longrightarrow\mathbb{R}/\mathcal{P},$
where $algCob^{or}(M,\omega)$ is the homologically trivial part of the oriented algebraic Lagrangian cobordism group of $(M,\omega)$, and $\mathcal{P}\subset\mathbb{R}$ is a finitely generated group of periods. These are given by integrating $\omega^{2}\wedge\beta$ over $(n+1)$-chains bounding homologically trivial Lagrangians, for any (n3)(n-3)-form $\beta$ on $M$. This is developed in Section 4.2. Cobordism groups $\mathcal{C}ob(M,\omega)$ carry a stronger version of these flux morphisms, see Lemma 3.4.
The equality from Equation (1.1) is then interpreted as an instance of a more general correspondence between tropical fluxes in a tropical torus $B$, and symplectic fluxes in the corresponding Lagrangian torus fibration $X(B)$. More concretely: if a nullhomologous tropical $k$-cycle $Z$ admits a Lagrangian lift $L_{Z}$, this correspondence gives $\Theta^{trop}(Z)=\Theta^{symp}(L_{Z})$ [^1] Note that this equality should be made much more precise. In reality, we define multiple maps $\Theta^{trop}$ and $\Theta^{symp}$, corresponding to different forms being integrated on the bounding chains. We should really say that given a particular tropical flux (i.e., a particular choice of tropical form on $B$), can be equated to a particular symplectic flux (i.e., a corresponding choice of (n3)(n-3)-form $\beta$) on $X(B)$..
In these more general terms, our proof of Theorem 1.1 simply consists of exhibiting an instance of the following commutative diagram:
$algCob^{or}(X(B))_{hom}$$\text{Griff}_{k}(B)_{hom}$$\mathbb{R}/\mathcal{P},$$\scriptstyle{symplectic\hskip 1.42262ptflux}$$\scriptstyle{Lagrangian\hskip 2.84526ptlift}$$\scriptstyle{tropical\hskip 0.56905ptflux}$
and concluding from Zharkov’s computation.
Remark 1.3.
Note that we show (Remark 4.5) that a Lagrangian cobordism between $L$ and $L^{\prime}$ gives rise to an algebraic Lagrangian cobordism between $L$ and $L^{\prime}$ (which is what we could expect from the fact that algebraic equivalence is weaker than rational equivalence). Therefore, an immediate Corollary of Theorem 1.1 is the following:
@@
For generic tropical curve of type $K4$, the Lagrangian Ceresa cycle $L_{C}-L_{C}^{-}$ has infinite order in $\mathcal{C}ob^{or}(X(J(C)))$.
Structure of the paper
Section 2 contains two parts. The first introduces the relevant background in tropical geometry before going through the construction of the tropical Ceresa cycle associated to a tropical curve. Special attention is given to the case of the $K4$ curve, which is of main interest to us.
The second part defines tropical flux morphisms using the language of tropical (co)homology. These are given by integrals of tropical forms, which are generalisations of Zharkov’s determinantal form.
In Section 3 we give definitions for the various flavours of Lagrangian cobordism groups. We introduce symplectic flux morphisms from $\mathcal{C}ob^{or}(M,\omega)$, and include an example application to the cotangent bundle of a torus.
Section 4 introduces the notion of algebraic Lagrangian cobordism. We then construct symplectic flux maps from algebraic Lagrangian cobordism groups $algCob(M,\omega)$. We end by showing that these fluxes can be encoded as a map to a Lefschetz Jacobian, a naturally tropical object which plays the role symplectically of intermediate Jacobians.
The aim of Section 5 is to formalise the relation between tropical fluxes in a tropical torus $B$, and symplectic fluxes in the corresponding Lagrangian torus fibration $X(B)$. Because both are given by integrating certain differential forms over chains, this involves constructing maps of (co)chains from tropical (co)chains in $B$ to singular (co)chains in $X(B)$ in a way that preserves the integration pairing. The underlying maps on (co)homology are a Künneth-type decomposition for $X(B)\cong B\times T^{n}$.
We prove our main Theorem 1.1 in Section 6 by bringing together elements of the previous Sections. This is done after an explicit description of the Lagrangian Ceresa cycle as a $3$-manifold inside the symplectic $6$-torus $X(J(C))$. We also speculate on a higher-dimensional extension of this result.
We end with two additional remarks in Section 7. The first turns to the case of genus $3$ tropical curves of hyperelliptic type, whose Ceresa cycles provide one with a rich source of interesting questions and have inspired recent work [ceresa_period_tropical, Ceresa_class_corey_et_al]. The second relates our work to a construction by Reznikov [Reznikov_characteristic_classes_in_symplectic_topology] building characters of symplectic Torelli-type groups.
Notation
We abuse notation and simply denote by $M$ a symplectic manifold $(M,\omega)$. Similarly, $(M,-\omega)$ will be written $\overline{M}$.
Acknowledgements
I am greatly endebted to Ivan Smith for his guidance and support, and for carefully reading previous drafts of this manuscript. I also thank an anonymous referee for providing constructive feedback which has improved the exposition of the paper. I was supported by EPSRC grant EP/X030660/1 for the duration of this work.
2. The tropical Ceresa cycle and tropical flux
This Section consists of two parts. The first starts by introducing the relevant background in tropical geometry, with a view to constructing the tropical Ceresa cycle associated to a tropical curve. It ends with a detailed study of the case which is of interest to us, namely that of the $K4$ curve.
The second introduces the notion of tropical flux using the language tropical (co)homology. These are morphisms from the homologically trivial part of the tropical Griffiths groups given by integrals of a generalised version of Zharkov’s determinantal form [zharkov_tropical_ceresa]. They formalise and generalise Zharkov’s argument for proving Theorem 1.2, where he exhibits some non-trivial integral over a chain bounding the tropical Ceresa cycle.
2.1. Tropical background
2.1.1. Tropical tori
Definition 2.1.
Remark 2.2.
The base $B$ of a complete integrable system $M\rightarrow B$ carries a canonical tropical affine structure given by the action-angle coordinates [johnnyevans_ltf, Section 2.3].
$0\longrightarrow\underline{\mathbb{R}}\longrightarrow\textit{Aff}_{B}\longrightarrow T_{\mathbb{Z}}^{*}B\longrightarrow 0.$
In the case where $B:=\mathbb{R}^{n}/\Gamma$ is a torus, the linear part of the monodromy on the sheaf of affine functions vanishes, and the local systems $T^{(*)}_{\mathbb{Z}}B$ are trivial. Therefore a tropical torus $B$ is simply the data of two rank $n$ lattices $\Gamma_{1}$ and $\Gamma_{2}$ in $\mathbb{R}^{n}$, such that $B=\mathbb{R}^{n}/\Gamma_{1}$ and the tropical structure on $B$ is given by $T_{\mathbb{Z}}B=\Gamma_{2}\subset\mathbb{R}^{n}$. Up to a global coordinate transformation one can take $\Gamma_{2}$ to be the standard lattice $\mathbb{Z}^{n}\subset\mathbb{R}^{n}$, then a tropical torus is determined by a matrix $Q$ in $GL(n;\mathbb{R})$ such that $B=B(Q):=\mathbb{R}^{n}/Q\cdot\mathbb{Z}^{n}$ (note that left or right multiplication of $Q$ by elements of $GL(n;\mathbb{Z})$ yield isomorphic tropical tori, see e.g. [nick_ivan1, Lemma 2.4]).
$H^{1}(\textit{Aff}_{B})\longrightarrow H^{1}(T_{\mathbb{Z}}^{*}B).$
$c:H_{1}(B;\mathbb{Z})\longrightarrow H^{0}(T_{\mathbb{Z}}^{*}B)$
2.1.2. Tropical curves, their Jacobian and their Ceresa cycle
Definition 2.3.
Remark 2.4.
As such, $C$ is not quite a tropical manifold of dimension $1$ as in Definition 2.1, but falls into the more general notion of tropical space [tropical_book, Definition 7.1.8]. The two can be shown to be equivalent [tropical_book, Section 8.1] if one additionally requires $C$ to be smooth, regular at infinity, and of finite type [tropical_book, Definition 7.4.1]. For our purposes these assumptions are not restrictive, and we extend our constructions on tropical manifolds to tropical curves, where the tropical structure on the latter can be obtained by embedding edges of $C$ in $\mathbb{R}$ via metric-preserving charts (for more details, consult e.g. [Mikhalkin_Zharkov_trop_curves_jac_theta_funct, Proposition 3.6]).
If the underlying graph of $C$ has genus $g$, then $H^{0}(T_{\mathbb{Z}}^{*}C)$ is a free abelian group of rank $g$; in particular $H^{0}(T_{\mathbb{R}}^{*}C):=H^{0}(T_{\mathbb{Z}}^{*}C)\otimes\mathbb{R}\cong\mathbb{R}^{g}$. We define the tropical Jacobian of $C$ as the torus
$J(C):=H^{0}(T_{\mathbb{R}}^{*}C)^{*}/H_{1}(C;\mathbb{Z}),$
where $H_{1}(C;\mathbb{Z})$ is identified with a lattice inside $H^{0}(T_{\mathbb{R}}^{*}C)^{*}$ by setting $\gamma(\alpha):=\int_{\gamma}\alpha$ for any $\alpha\in H^{0}(T_{\mathbb{R}}^{*}C)$.
Given a choice of basepoint $p\in C$, the tropical Abel-Jacobi map
$AJ_{p}:\hskip 5.69054pt$ $C\longrightarrow J(C)$
$Q(E,E):=l(E)$
on any edge $E$ of $C$ can be extended bilinearly. This induces a symmetric, positive-definite bilinear form on $H_{1}(C;\mathbb{Z})\cong H_{1}(J(C);\mathbb{Z})$, therefore a polarisation
$c(C):H_{1}(J(C);\mathbb{Z})\longrightarrow H^{0}(T^{*}_{\mathbb{Z}}J(C))$
by $\int_{\gamma_{i}}c(\gamma_{j}):=Q(\gamma_{i},\gamma_{j})$.
$(-1):J(C)\longrightarrow J(C)$
given by inverse for the group law. The Ceresa cycle of $C$ is
$AJ_{p}(C)-(-1)_{*}(AJ_{p}(C))\in CH_{1}(J(C)).$
2.1.3. An important example: the genus $3$ curve of type $K4$
Here we explore the case where $C$ is a genus $3$ curve of type $K4$ represented in Figure 2; which is of primary interest to us. It is one of five types of generic genus $3$ tropical curves, and distinguishes itself by being the only non-hyperelliptic type. A tropical curve is of hyperelliptic type if its Jacobian is isomorphic to the Jacobian of a hyperelliptic tropical curve.
ppEFDACB
Figure 2. A tropical curve of type $K4$.
When a curve is hyperelliptic, its Ceresa cycle is trivial: this can be seen by embedding it in its Jacobian using as a basepoint a Weierstrass point $p\in C$. Then one finds that $AJ_{p}(C)=(-1)_{*}(AJ_{p}(C))$. However, as soon as $g(C)\geq 3$, the Ceresa cycle (whether in its classical or tropical version) has been the object of extensive study for being an example of a cycle which is homologically, but not algebraically, trivial. The following result is originally due to Ceresa in the classical setting [Ceresa], but its tropical version was proved by Zharkov: